Calculate Ecell at 298 K for a cell involving Sn and Cu and their ions:

Sn(s) I Sn2+(aq, 0.25M) II Cu2=(aq, 0.10M) I Cu(s)

A) 0.47 V

B) 0.49 V

C) 0.50 V
Answer #1

Eo (Sn2+/Sn) = -0.14 V
Eo (Cu2+/Cu) = 0.34 V

Eo cell = Eo right - Eo left
= 0.34 - (-0.14)
= 0.48 V

the reaction taking place is:
Sn + Cu2+ ----> Sn2+ + Cu
number of electron, n = 2
use:
Ecell = Eocell - 0.059/2 * log ([Sn2+]/[Cu2+])
= 0.48 - 0.059/2 * log (0.25/0.10)
= 0.48 - 0.059/2 * log (2.5)
= 0.468 V
= 0.47 V (approx)
Answer: A